OpenStudy (anonymous):

Find the value of x if the area of the triangle is 357 in2.

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

area of triangle = 1/2(bh) 357in^2 = 1/2(34x)

OpenStudy (anonymous):

357in^2=17x?

OpenStudy (anonymous):

is tht the next step

OpenStudy (aripotta):

357in^2 = 1/2(34x) you'll wanna multiply both sides by 2 714in^2 = 34x

OpenStudy (anonymous):

ohh and evide by 34 to both sides?

OpenStudy (aripotta):

yea

OpenStudy (anonymous):

thx :D

OpenStudy (aripotta):

no problemo